OpenStudy (anonymous):

An investment pays simple interest, and doubles in 7 years. What is the interest rate? Simple interest = A=P(1+rt)

OpenStudy (jdoe0001):

looks like you just need to solve for "r"

OpenStudy (anonymous):

What is my principle?

OpenStudy (jdoe0001):

well. I guess is P and the balance after 7 years is A

OpenStudy (jdoe0001):

I mean... we don't have constant value for those

OpenStudy (anonymous):

With so little information how would I figure the interest rate? I am not familiar with the steps.

OpenStudy (jdoe0001):

well years is "t" and is pretty much all that's given

OpenStudy (anonymous):

so do I put in any number?

OpenStudy (jdoe0001):

so A=P(1+rt) => A=P(1+r(7))

OpenStudy (jdoe0001):

I'd think not, I'd think you'd solve for "r" in A and P terms

OpenStudy (anonymous):

A=P(1+r(7)) 2=1(1+r(7)) How do I solve?

OpenStudy (jdoe0001):

2? ... we dunno what A is so is just A=P(1+r(7)) distribute the P and just do a linear equation simplification

OpenStudy (anonymous):

I think I got it.

OpenStudy (anonymous):

2=1(1+r(7) 2=(1+r7) 2-1=(1+r7)-1 1/7=r7/7 0.14285=r thus the interest rate is 14.3%

OpenStudy (jdoe0001):

hmmm I never did see the "2" or the "1"

OpenStudy (anonymous):

I think you just take a number that doubles and it will always be 1.

OpenStudy (anonymous):

I mean 1/7

OpenStudy (jdoe0001):

hmmm

OpenStudy (jdoe0001):

yeap.... you're right...

OpenStudy (anonymous):

woohoo Thanks for helping. I don't know what I'd do without this site.

OpenStudy (jdoe0001):

yw